OpenStudy (anonymous):

The antenna tank of a Graymark 536 Radio Kit has an inductance L≈0.65mH. The resistance of the inductor RL≈4.0Ω. The equivalent resistance across the capacitor is RC≈490.0kΩ. The capacitor is variable, for tuning. If we tune to a station at f=920.0kHz what is the capacitance, in picoFarads, of the tuning capacitor? please give me the formula ????

OpenStudy (anonymous):

H11P1 a) 1/(RC*C) + RL/L b) (thats the formula = 1/(2*pi*f)^(2)*L) c) range of dis ans within 4 and 10 (ALSO USE DECIMAL VALUES) d) (thats the formula = 1/(2*pi*f)^(2)*L) e) range of dis ans is from 15 and 20 (ALSO USE DECIMAL VALUES)

OpenStudy (anonymous):

the range in (e) isn't fro 15 to 20. I've got 21 !

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)